# Extreme Conformal Prediction: Reliable Intervals for High-Impact Events

### v1

@@ Abstract
Conformal prediction is a popular method to construct prediction intervals for black-box machine learning models with marginal coverage guarantees. In applications with potentially high-impact events, such as flooding or financial crises, regulators often require very high confidence for such intervals. However, if the desired level of confidence is too large relative to the amount of data used for calibration, then classical conformal methods provide infinitely wide, thus, uninformative prediction intervals. In this paper, we propose a new method to overcome this limitation. We bridge extreme value statistics and conformal prediction to provide reliable and informative prediction intervals with high-confidence coverage, which can be constructed using any black-box extreme quantile regression method. The advantages of this extreme conformal prediction method are illustrated in a simulation study and in an application to flood risk forecasting.
@@ 1 Introduction
Conformal prediction is a simple approach to producing prediction sets from any regression or classification model. For a covariate vector $\bm{X}$ with values in $\mathcal{X}\subseteq\mathbb{R}^{p}$ and corresponding response variable $Y$, the goal of classical conformal prediction is to build a prediction set $C(\mathbf{x})$ satisfying marginal coverage
$\mathbb{P}(Y_{\text{test}}\in C(\bm{X}_{\text{test}}))\geq 1-\alpha,$ (1)
@@
for a desired confidence level $1-\alpha\in(0,1)$, for any new observations $(\bm{X}_{\text{test}},Y_{\text{test}})$. To obtain such a prediction set, we assume that a prediction model $\hat{f}$ was fitted on a training data set from the distribution $\mathcal{P}$ of $(\mathbf{X},Y)$, and that a new calibration data set $(\mathbf{X}_{1},Y_{1}),\dots,(\mathbf{X}_{n_{c}},Y_{n_{c}})$ of $n_{c}$ independent samples from the same distribution is available. For a specific score function $s:\mathcal{X}\times\mathbb{R}\to\mathbb{R}$ that may depend on $\hat{f}$ and acts on the predictors and responses, consider the calibration scores $s_{i}=s(\mathbf{X}_{i},Y_{i})$ for $i=1,\dots,n_{c}$. For some level $\alpha\in(0,1)$, denoting by $\hat{q}_{\alpha}$ the $(\left\lceil(n_{c}+1)(1-\alpha)\right\rceil/n_{c})$-quantile of the scores $s_{1},\dots,s_{n_{c}}$, the prediction set
$C(\mathbf{x})=\{y:s(\mathbf{x},y)\leq\hat{q}_{\alpha}\}.$ (2)
@@
has the desired $1-\alpha$ coverage. That is, it satisfies Eq. 1 for a new test sample $(\bm{X}_{\text{test}},Y_{\text{test}})$ from the same distribution $\mathcal{P}$. This well-established framework is the so-called split conformal approach (e.g. Lei et al., 2018; Romano et al., 2019). More generally, originally started in Vovk et al. (2022), the conformalization idea that leverages quantile-based construction of prediction sets elicits a range of variants, with focus on optimal data usage and applying to different problems, including jackknife+ (Alaa and van der Schaar, 2020; Barber et al., 2021), cross-conformal prediction (Vovk, 2015) and ensemble-based approaches (Kim et al., 2020; Gupta et al., 2022).
@@
Conformal prediction intervals are widely used for confidence levels $1-\alpha$ of moderate value relative to the sample size $n_{c}$ (e.g., $90\%$ and $n_{c}=1000$), where enough of the calibration scores are above this quantile. In many applications, however, test points $Y_{\text{test}}\notin C(\bm{X}_{\text{test}})$ that fall outside of the confidence set correspond to a high-impact event with serious consequences for the environment, human lives or the economy. Examples for such risk-sensitive applications are the protection of cities and energy infrastructure from flooding (Keef et al., 2009; Asadi et al., 2015) or the financial reserves of banks and insurance companies (van Oordt and Zhou, 2019; Dupuis et al., 2023). In these cases, much larger values of confidence $1-\alpha$, close to one, will be required, sometimes even by law. Classical methods from conformal prediction fail for those requirements, since the quantile $\hat{q}_{\alpha}$ as defined above is not a useful estimator when the level $\alpha<1/(n_{c}+1)$. Indeed, in this case, less than one observation then exceeds the $(1-\alpha)$-quantile on average, and $\hat{q}_{\alpha}$ is infinite (or ill-defined). Even for slightly larger values $\alpha$ close to that limit, the variance of $\hat{q}_{\alpha}$ can be huge.
@@
Extreme value theory provides statistical tools for accurate estimation beyond the data range (de Haan and Ferreira, 2006). The tools have proven successful for improving extrapolation properties of machine learning methods in regression (de Carvalho et al., 2022; Huet et al., 2024; Buriticá and Engelke, 2024), classification (Jalalzai et al., 2018) and generative methods (Boulaguiem et al., 2022).
@@
In this paper, we propose a new methodology that bridges the wide applicability of conformal prediction with extrapolation tools from extreme value statistics to construct reliable prediction sets for high-impact events. In a first step, in order to obtain a good pretrained model $\hat{f}$ beyond the data range, we rely on flexible machine learning methods from extreme quantile regression (Velthoen et al., 2019; Gnecco et al., 2024; Pasche and Engelke, 2024; Richards and Huser, 2024). Second, we rely on the classical and theoretically justified peaks-over-threshold approach, which consists of using the generalised Pareto distribution (GPD) to extrapolate, for example, quantile estimates beyond the range of empirical observations (Balkema and de Haan, 1974; Pickands III, 1975). For a confidence level $1-\alpha$ close to one, we leverage the GPD fitted to the calibration scores $s_{1},\dots,s_{n_{c}}$ to obtain a reliable estimate of $\hat{q}_{\alpha}$ beyond the calibration data. The resulting extreme conformal prediction intervals have better properties compared to those from the classical empirical approach for large confidence requirements. In a simulation study, we show that our method improves existing approaches in terms of better coverage, in the sense of Eq. 1, and of informativeness of the prediction interval.
@@
We illustrate the advantages of our approach in an application to flood risk assessment. Using several of the flexible machine learning methods as base predictions, it provides high-confidence one-day-ahead interval forecasts of the conditional range for water flow. We show that using conformal prediction intervals based on extreme value theory improves the coverage of the classical method, which either yields uninformative intervals or exhibits undercoverage and, therefore, seriously underestimates the risk of high-impact events.
@@ 2.1 Split conformal prediction
As in the introduction, let $\bm{X}$ be a covariate vector taking values in $\mathcal{X}\subseteq\mathbb{R}^{p}$, and $Y$ the response random variable of interest. We will consider, here and in the sequel, the regression case where the response $Y$ is real-valued in $\mathbb{R}$. We also suppose that we have access to a calibration set of $n_{c}$ independent observations $(\mathbf{X}_{1},Y_{1}),\dots,(\mathbf{X}_{n_{c}},Y_{n_{c}})$. Classical (split) conformal prediction builds prediction intervals (PIs) as in Eq. 2 that have desired coverage under very weak assumptions. For instance, if the calibration set is an exchangeable sample, and the test point $(\bm{X}_{\text{test}},Y_{\text{test}})$ is independent and of the same distribution, then the unconditional coverage guarantee Eq. 1 holds (Papadopoulos et al., 2002).
Importantly, the probability measure in Eq. 1 is with respect to the randomness in the calibration set jointly with the test point, that is, $\{(\bm{X}_{1},Y_{1}),\ldots,(\bm{X}_{n_{c}},Y_{n_{c}}),(\bm{X}_{\text{test}},Y_{\text{test}})\}$. In fact, when conditioning on the calibration set, the distribution of the coverage is a beta distribution, that is,
$\mathbb{P}\left(Y_{\text{test}}\in C(\bm{x})\mid\{(\bm{X}_{i},Y_{i})\}_{i=1}^{n_{c}}\right)\sim{\rm Beta}(n+1-l,\,l),\quad l:=\left\lfloor(n+1)\alpha\right\rfloor.$ (3)
Running the conformal prediction twice on different calibration sets, therefore yields PIs with different coverage probabilities. The guarantee in Eq. 1 says that when averaging out the calibration set, the coverage is at least $1-\alpha$.
Furthermore, the marginal coverage property in Eq. 1 only guarantees “overall” marginal coverage of the prediction set $C(\bm{x})$, but does not imply the conditional coverage property
$\mathbb{P}(Y_{\text{test}}\in C(\bm{x})\mid\bm{X}_{\text{test}}=\bm{x})\geq 1-\alpha,\quad\forall\bm{x}\in\mathcal{X}.$ (4)
The latter is generally impossible to guarantee in such a general setting. How close $C(\bm{x})$ is to satisfying the conditional coverage property depends on the quality of the given pretrained model. For example, for the conformalized quantile regression approach described in Section 2.2, it depends on the accuracy of the initial quantile regression model $\hat{Q}_{\alpha/2}(\bm{x}),\hat{Q}_{1-\alpha/2}(\bm{x})$.
@@ 2.2 Conformalized quantile regression
Conformalized quantile regression was first proposed by Romano et al. (2019) and is also described in Angelopoulos and Bates (2023). Suppose that we have access to a black-box quantile regression model trained to estimate the conditional quantiles $\hat{Q}_{\alpha/2}(\bm{x})$ and $\hat{Q}_{1-\alpha/2}(\bm{x})$ of $Y$, given $\bm{X}=\bm{x}$, at probability levels $\alpha/2$ and $21-\alpha/$ 2, respectively. Then, conformalized quantile regression uses the score function
@@
Following the general procedure described in the introduction, this leads to the final prediction set in Eq. 2 being the interval
$C(\bm{x})=\left[\hat{Q}_{\alpha/2}(\bm{x})-\hat{q}_{\alpha},\hat{Q}_{1-\alpha/2}(\bm{x})+\hat{q}_{\alpha}\right],$ (6)
where $\hat{q}_{\alpha}$ is the $(\left\lceil(n_{c}+1)(1-\alpha)\right\rceil/n_{c})$-quantile of the calibration scores $s_{1},\ldots,s_{n_{c}}$, i.e. the order statistic $s_{(\left\lceil(n_{c}+1)(1-\alpha)\right\rceil)}$.
Intuitively, the procedure either widens (with a positive $\hat{q}_{\alpha}$) or narrows (with a negative $\hat{q}_{\alpha}$) the initial interval $\left[\hat{Q}_{\alpha/2}(\bm{x}),\hat{Q}_{1-\alpha/2}(\bm{x})\right]$ so that it covers $\left\lceil(n_{c}+1)(1-\alpha)\right\rceil$ of the $n_{c}$ calibration observations. Note that the resulting prediction intervals satisfy the marginal coverage Eq. 1, but there is no guarantee that the conditional coverage in Eq. 4 is satisfied. In fact, the more accurate the initial quantile regression models $\hat{Q}_{\alpha/2}(\bm{x})$ and $\hat{Q}_{1-\alpha/2}(\bm{x})$ are, the better the conditional coverage will be.
@@ 2.3 Limitation for extreme confidence levels
For high-impact events, regulators often require predictions with very high coverage probabilities to ensure that protective infrastructures or measures are sufficient. In particular, in such risk-sensitive applications, the level $\alpha$ in Eq. 1 is typically close to 00 and may satisfy $\alpha<1/(n_{c}+1)$. This is generally referred to as an extreme confidence or probability level since, on average, there is less than one observation above the $(1-\alpha)$-quantile in a sample of size $n_{c}$. Note that the size $n_{c}$ of the calibration set is typically fairly small, since these data cannot be used for model fitting, often resulting in extreme scenarios even for relatively moderate levels of $\alpha$.
The classical construction of the conformal prediction intervals described in the introduction requires the computation of $\hat{q}_{\alpha}$, the empirical $(\left\lceil(n_{c}+1)(1-\alpha)\right\rceil/n_{c})$-quantile of the calibration set scores. For extreme confidence levels $1-\alpha$, this quantile level
@@
in which case $\hat{q}_{\alpha}$ is ill-defined and, by convention, set to infinity (Romano et al., 2019; Angelopoulos and Bates, 2023). This results in degenerate trivial prediction intervals $C(\bm{x})=\left(-\infty,\infty\right)$, for all $\bm{x}\in\mathcal{X}$. Although this interval satisfies the coverage Eq. 1, it is of no practical utility.
@@ 3 Extreme conformal prediction
We propose an approach based on extreme value statistics to construct non-degenerate conformal prediction intervals at extreme confidence levels $1-\alpha>n_{c}/(n_{c}+1)$. Similarly to classical conformalized quantile regression (Romano et al., 2019), our method requires two steps:
@@
    calibrating the scores $s_{1},\dots,s_{n_{c}}$ on an independent data set.
For extreme confidence levels, both steps typically require extrapolation beyond the data range. Indeed, if $\alpha$ is close to 0, and in particular if $\alpha<1/(n+1)$ is also extreme in the training data, then usual quantile regression will not be accurate. Instead, extreme quantile regression methods should be used. There is large literature on such methods based on linear models (Chernozhukov, 2005), additive models (Chavez-Demoulin and Davison, 2005; Youngman, 2019), or more flexible machine learning models such as gradient boosting (Velthoen et al., 2023), random forest (Gnecco et al., 2024) or neural networks (Pasche and Engelke, 2024; Richards and Huser, 2024). Importantly, the model in step 1. can be a black-box, in the sense that we do not require theoretical guarantees. We discuss the extrapolation for step 2. in this section and come back to examples of extreme quantile regression models in Sections 4 and 5.
3.1 Unilateral prediction intervals
Extreme conformal prediction intervals are most relevant in cases where very large values of the response variable $Y$ lead to severe negative impacts. In such cases, reliable prediction intervals for $Y\mid\bm{X}=\bm{x}$, which contain the realisation of $Y$ with very high marginal probability, are a crucial forecasting tool. They can be used to determine whether a dangerous level of the response could potentially be reached. This also allows the reduction of false negatives (e.g., in the form of missing warnings) that can be critical to the system. In those risk assessment scenarios, it is often a single of the two tail directions which is of risk. This is the case for the risk of flooding discussed in Section 5 but also, for instance, for high temperatures in dry areas at risk of wildfires, and financial asset returns at risk of large losses. Without loss of generality, we, thus, suppose that one is interested in single-sided prediction intervals; two-sided intervals can be constructed analogously.
The classical bilateral procedure can be adapted to unilateral prediction intervals using the score function

### v2

@@ Abstract
Conformal prediction is a popular method to construct prediction intervals for black-box machine learning models with marginal coverage guarantees. In applications with potentially high-impact events, such as flooding or financial crises, regulators often require very high confidence for such intervals. However, if the desired level of confidence is too large relative to the amount of data used for calibration, then classical conformal methods provide infinitely wide, thus, uninformative prediction intervals. In this paper, we propose a new method to overcome this limitation. We bridge extreme value statistics and conformal prediction to provide reliable and informative prediction intervals with high-confidence coverage, which can be constructed using any black-box extreme quantile regression method. A weighted version of our approach can account for nonstationary data. The advantages of our extreme conformal prediction method are illustrated in a simulation study and in an application to flood risk forecasting.
@@ 1 Introduction
Conformal prediction is a simple approach to producing prediction sets from any regression or classification model. For a covariate vector $\bm{X}$ with values in $\mathcal{X}\subseteq\mathbb{R}^{p}$ and corresponding response variable $Y$, the goal of classical conformal prediction is to build a prediction set $\hat{C}(\bm{x})$ satisfying marginal coverage
$\mathbb{P}\{Y_{\text{test}}\in\hat{C}(\bm{X}_{\text{test}})\}\geq 1-\alpha,$ (1)
@@
for a desired confidence level $1-\alpha\in(0,1)$, for any new observations $(\bm{X}_{\text{test}},Y_{\text{test}})$. To obtain such a prediction set, we assume that a prediction model $\hat{f}$ was fitted on a training data set from the distribution $\mathcal{P}$ of $(\bm{X},Y)$, and that a new calibration data set $(\bm{X}_{1},Y_{1}),\dots,(\bm{X}_{n_{c}},Y_{n_{c}})$ of $n_{c}$ observations from the same distribution is available. For a specific nonconformity score function $s:\mathcal{X}\times\mathbb{R}\to\mathbb{R}$ that may depend on $\hat{f}$ and acts on the predictors and responses, consider the calibration scores $S_{i}=s(\bm{X}_{i},Y_{i})$ for $i=1,\dots,n_{c}$. For some level $\alpha\in(0,1)$, denoting by $\hat{q}_{\alpha}$ the $\{\left\lceil(n_{c}+1)(1-\alpha)\right\rceil/n_{c}\}$-quantile of the scores $S_{1},\dots,S_{n_{c}}$, the prediction set
$\hat{C}(\bm{x})=\{y:s(\bm{x},y)\leq\hat{q}_{\alpha}\}.$ (2)
@@
has the desired $1-\alpha$ coverage. That is, it satisfies Eq.˜1 for any new test observation $(\bm{X}_{\text{test}},Y_{\text{test}})$ from the same distribution $\mathcal{P}$, under a remarkably weak exchangeability assumption. This well-established framework is the so-called split conformal approach (SplitConformalMargProof; Lei18ConformalPred). More generally, originally started in vovk1999ml; vovk2005algorithmic, the conformalization idea that leverages quantile-based construction of prediction sets elicits a range of variants, with focus on optimal data usage and applying to different problems, including jackknife+ (alaa2020discriminative; barber2021predictive), cross-conformal prediction (vovk2015cross) and ensemble-based approaches (kim2020predictive; gupta2022nested). In combination with machine learning methods, it can efficiently capture predictive uncertainty (shafervovk2008tutorial; Zhou25ConformSurveyML), and provide intervals with highly adaptive lengths (ConformalizedQR). Extensions to nonexchangeable data are also well-studied (OliveiraConformNonexchError). In particular, weighted conformal methods can account for distribution shifts and drifts (TibshiraniCoformShift; Barber2023nonstatConform).
@@
Conformal prediction intervals are widely used for confidence levels $1-\alpha$ of moderate value relative to the sample size $n_{c}$ (e.g., $90\%$ and $n_{c}=1000$), where enough of the calibration scores are above this quantile. In many applications, however, test points $Y_{\text{test}}\notin\hat{C}(\bm{X}_{\text{test}})$ that fall outside of the confidence set correspond to a high-impact event with serious consequences for the environment, human lives or the economy. Examples for such risk-sensitive applications are the protection of cities and energy infrastructure from flooding (kee2009; AsadiDanube) or the financial reserves of banks and insurance companies (zhou2019; dup2022). In these cases, much larger values of confidence $1-\alpha$, close to one, will be required, sometimes even by law. Classical methods from conformal prediction fail for those requirements, since the quantile $\hat{q}_{\alpha}$ as defined above is not a useful estimator when the level $\alpha<1/(n_{c}+1)$. Indeed, in this case, less than one observation then exceeds the $(1-\alpha)$-quantile on average, and $\hat{q}_{\alpha}$ is infinite (or ill-defined). Even for slightly larger values $\alpha$ close to that limit, the variance of $\hat{q}_{\alpha}$ can be huge.
@@
Extreme value theory provides statistical tools for accurate estimation beyond the data range (DF2006). The tools have proven successful for improving extrapolation properties of machine learning methods in regression (de2022regression; huet2024regressionextremeregions; bur2024), classification (jalalzai:clemencon:sabourin:2018), and generative methods (bou2022).
@@
In this paper, we propose a new methodology that bridges the wide applicability of conformal prediction with extrapolation tools from extreme value statistics to construct reliable prediction sets for high-impact events. In a first step, in order to obtain a good pretrained model $\hat{f}$ beyond the data range, we rely on flexible machine learning methods from extreme quantile regression (Velthoenetal2019; erf; Pasche2024; RichardsNNfire). Second, we rely on the classical and theoretically justified peaks-over-threshold approach, which consists of using the generalised Pareto distribution (GPD) to extrapolate, for example, quantile estimates beyond the range of empirical observations (GPDBalkema; GPDPickands). For a confidence level $1-\alpha$ close to one, we leverage the GPD fitted to the calibration scores $S_{1},\dots,S_{n_{c}}$ to obtain a reliable estimate of $\hat{q}_{\alpha}$ beyond the calibration data. The resulting extreme conformal prediction intervals have better properties compared to those from the classical empirical approach for large confidence requirements. In a simulation study, we show that our method improves existing approaches in terms of better coverage, in the sense of Eq.˜1, and of informativeness of the prediction interval. We also consider a weighted version of our extreme conformal approach to deal with nonexchangeable data, and discuss its usage with several classical types of conformal procedures.
@@
The advantages of our approach are illustrated in an application to flood risk prediction. Using several of the flexible machine learning methods as base predictions, it provides high-confidence one-day-ahead interval forecasts of the conditional range for water flow. We show that using conformal prediction intervals based on extreme value theory improves the coverage of the classical method, which either yields uninformative intervals or exhibits undercoverage and, therefore, seriously underestimates the risk of high-impact events.
@@ 2.1 Split conformal prediction
As in the introduction, let $\bm{X}$ be a covariate vector taking values in $\mathcal{X}\subseteq\mathbb{R}^{p}$, and $Y$ the response random variable of interest. We will consider, here and in the sequel, the regression case where the response $Y$ is real-valued in $\mathbb{R}$. We also suppose that we have access to a calibration set of $n_{c}$ observations $(\bm{X}_{1},Y_{1}),\dots,(\bm{X}_{n_{c}},Y_{n_{c}})$. Classical split conformal prediction builds prediction intervals (PIs) as in Eq.˜2 that have desired coverage under very weak assumptions. Specifically, if the joint distribution of the calibration set and the test point $(\bm{X}_{\text{test}},Y_{\text{test}})$ is exchangeable, then the unconditional coverage guarantee Eq.˜1 holds (SplitConformalMargProof).
Importantly, the probability measure in Eq.˜1 is with respect to the randomness in the calibration set jointly with the test point, that is, $\{(\bm{X}_{1},Y_{1}),\ldots,(\bm{X}_{n_{c}},Y_{n_{c}}),(\bm{X}_{\text{test}},Y_{\text{test}})\}$. In fact, when conditioning on the calibration set, the distribution of the coverage is a beta distribution, that is,
$\mathbb{P}\left[Y_{\text{test}}\in\hat{C}(\bm{X}_{\text{test}})\mid\{(\bm{X}_{i},Y_{i})\}_{i=1}^{n_{c}}\right]\sim{\rm Beta}(n_{c}+1-l,\,l),\quad l:=\left\lfloor(n_{c}+1)\alpha\right\rfloor.$ (3)
Running the conformal prediction twice on different calibration sets, therefore yields PIs with different coverage probabilities. The guarantee in Eq.˜1 says that when averaging out the calibration set, the coverage is at least $1-\alpha$.
Furthermore, the marginal coverage property in Eq.˜1 only guarantees “overall” marginal coverage of the prediction set $\hat{C}(\bm{x})$, but does not imply the conditional coverage property
$\mathbb{P}\{Y_{\text{test}}\in\hat{C}(\bm{x})\mid\bm{X}_{\text{test}}=\bm{x}\}\geq 1-\alpha,\quad\forall\bm{x}\in\mathcal{X}.$ (4)
The latter is generally impossible to guarantee in such a general setting. How close $\hat{C}(\bm{x})$ is to satisfying the conditional coverage property depends on the quality of the given pretrained model. For example, for the conformalized quantile regression approach described in Section˜2.2, it depends on the accuracy of the initial quantile regression model $\hat{Q}_{\alpha/2}(\bm{x}),\hat{Q}_{1-\alpha/2}(\bm{x})$.
@@ 2.2 Conformalized quantile regression
Conformalized quantile regression, first proposed by ConformalizedQR and also described in ConformalIntrRev, is one of the most popular conformal methods, in particular thanks to its ability to provide varying-length prediction intervals with competitive adaptivity. Suppose that we have access to a black-box quantile regression model trained to estimate the conditional quantiles $\hat{Q}_{\alpha/2}(\bm{x})$ and $\hat{Q}_{1-\alpha/2}(\bm{x})$ of $Y$, given $\bm{X}=\bm{x}$, at probability levels $\alpha/2$ and $1-\alpha/2$, respectively. Then, conformalized quantile regression uses the score function
@@
Following the general procedure described in the introduction, this leads to the final prediction set in Eq.˜2 being the interval
$\hat{C}(\bm{x})=\left[\hat{Q}_{\alpha/2}(\bm{x})-\hat{q}_{\alpha},\;\hat{Q}_{1-\alpha/2}(\bm{x})+\hat{q}_{\alpha}\right],$ (6)
where $\hat{q}_{\alpha}$ is the empirical $\{\left\lceil(n_{c}+1)(1-\alpha)\right\rceil/n_{c}\}$-quantile of the calibration scores $S_{1},\ldots,S_{n_{c}}$, i.e. the order statistic $S_{(\left\lceil(n_{c}+1)(1-\alpha)\right\rceil)}$. As an equivalent definition, $\hat{q}_{\alpha}$ equals the empirical $(1-\alpha)$-quantile of $\{S_{1},\ldots,S_{n_{c}}\}\cup\{+\infty\}$, the calibration score sample augmented with a point at infinity.
Intuitively, the procedure either widens (with a positive $\hat{q}_{\alpha}$) or narrows (with a negative $\hat{q}_{\alpha}$) the initial interval $\left[\hat{Q}_{\alpha/2}(\bm{x}),\hat{Q}_{1-\alpha/2}(\bm{x})\right]$ so that it covers $\left\lceil(n_{c}+1)(1-\alpha)\right\rceil$ of the $n_{c}$ calibration observations. Note that the resulting prediction intervals satisfy the marginal coverage Eq.˜1, but there is no guarantee that the conditional coverage in Eq.˜4 is satisfied. In fact, the more accurate the initial quantile regression models $\hat{Q}_{\alpha/2}(\bm{x})$ and $\hat{Q}_{1-\alpha/2}(\bm{x})$ are, the better the conditional coverage will be.
@@ 2.3 Limitation for extreme confidence levels
For high-impact events, regulators often require predictions with very high coverage probabilities to ensure that protective infrastructures or measures are sufficient. In particular, in such risk-sensitive applications, the level $\alpha$ in Eq.˜1 is typically close to 0 and may satisfy $\alpha<1/(n_{c}+1)$. This is generally referred to as an extreme confidence or probability level since, on average, there is less than one observation above the $(1-\alpha)$-quantile in a sample of size $n_{c}$. Note that the size $n_{c}$ of the calibration set is typically fairly small, since these data cannot be used for model fitting, often resulting in extreme scenarios even for relatively moderate levels of $\alpha$.
The classical construction of the conformal prediction intervals described in the introduction requires the computation of $\hat{q}_{\alpha}$, the empirical $\{\left\lceil(n_{c}+1)(1-\alpha)\right\rceil/n_{c}\}$-quantile of the calibration set scores. For extreme confidence levels $1-\alpha$, this quantile level
@@
in which case $\hat{q}_{\alpha}$ is ill-defined and, by convention, set to infinity (ConformalizedQR; ConformalIntrRev). This results in degenerate trivial prediction intervals $\hat{C}(\bm{x})=\left(-\infty,\infty\right)$, for all $\bm{x}\in\mathcal{X}$. Although this interval satisfies the coverage Eq.˜1, it is of no practical utility.
@@ 3 Extreme conformal prediction
We propose an approach based on extreme value statistics to construct nondegenerate conformal prediction intervals at extreme confidence levels $1-\alpha>n_{c}/(n_{c}+1)$. Similarly to classical conformalized quantile regression (ConformalizedQR), our method requires two steps:
@@
    calibrating based on the scores $S_{1},\dots,S_{n_{c}}$ on an independent data set.
For extreme confidence levels, both steps typically require extrapolation beyond the data range. Indeed, if $\alpha$ is close to 0, and in particular if $\alpha<1/(n+1)$ is also extreme in the training data, then usual quantile regression will not be accurate. Instead, extreme quantile regression methods should be used. There is large literature on such methods based on linear models (chernozhukov2005), additive models (ExGAM; ExGAM2; de2022bayesianLRtails), or more flexible machine learning models such as gradient boosting (gbex; Kohfireboosting), random forest (erf) or neural networks (Pasche2024; RichardsNNfire; AlloucheNN). Importantly, the model in step 1 can be a black-box, in the sense that we do not require theoretical guarantees. We discuss the extrapolation for step 2 in Section˜3.2 and come back to examples of extreme quantile regression models in Sections˜4 and 5. Furthermore, we discuss extensions of our approach to nonexchangeable data and alternative conformal procedures in Sections˜3.3 and A.1.
3.1 Single-sided prediction intervals
Extreme conformal prediction intervals are most relevant in cases where very large values of the response variable $Y$ lead to severe negative impacts. In such cases, reliable prediction intervals for $Y\mid\bm{X}=\bm{x}$, which contain the realisation of $Y$ with very high marginal probability, are a crucial forecasting tool. They can be used to determine whether a dangerous level of the response could potentially be reached. This also allows the reduction of false negatives (e.g., in the form of missing warnings) that can be critical to the system. In those risk assessment scenarios, it is often a single of the two tail directions which is of risk. This is the case for the risk of flooding discussed in Section˜5 but also, for instance, for high temperatures in dry areas at risk of wildfires, and financial asset returns at risk of large losses. Without loss of generality, we, thus, suppose that one is interested in single-sided prediction intervals; two-sided intervals can be constructed analogously.
The classical procedure, yielding two-sided intervals, can be adapted to obtain single-sided prediction intervals by using the score function